# Consistent Validation for Predictive Methods in Spatial Settings

### v2

@@ E.6.2 Loss Function
E.6.3 Estimation of Ground Truth
Following the argument in Section E.1, we use the empirical test risk Equation 112 in place of the test risk as ground truth in synthetic experiments. The assumption that $(Z_{m})_{m=1}^{M^{\textup{{test}}}}$ are independent. It is sufficient for $Y^{\textup{{test}}}_{m}$ (the wind daily wind speeds) to be independent, conditioned on the spatial location at which it is observed. This is likely not the case, as we would expect average wind speed in consecutive days exhibit at least some dependence. However, if the wind speed decorrelates reasonably rapidly over time, we would expect similar arguments to hold, possibly with fewer effective samples.
Because the loss is truncated mean squared error, the $Z_{m}$ are surely bounded by $25$m2/s2. We use $775$ samples in estimating the risk. Using these numbers, and under the assumption that wind speed at a location is independent of the wind speed on previous days, in Equation 115, we arrive at
@@ E.6.3 Estimation of Ground Truth
Comparing to Figure 12, we see that this application of Hoeffding’s inequality is not sufficient to justify that the estimate of ground truth is accurate enough to attributed the observed better performance of SNN to (actually) better estimation of the ground truth as opposed to inaccuracies of our Monte Carlo estimate of the test risk. However, we expect this is largely due to looseness is Hoeffding’s inequality and, given that a large difference is observed in most seeds, it would be very surprising if this was only due to error in estimation of the ground truth which is independent across seeds.
E.6.4 Model Fitting
A gradient boosted machine is fit using LightGBM [Ke et al., 2017] with default parameters expect for the number of leaves (set to 127) and the number of estimators (set to 100).
E.6.5 Risk Estimation procedures
Holdout, 1NN and SNN are run as in the air temperature experiments (Section E.4.5). In particular, we use a failure probability of $\delta=0.1$ for SNN and nearest neighbor calculations are done with respect to Haversine distance to account for Earth’s curvature. We use a fixed $\delta$ since we only discuss a method applicable to estimating fill distance on the unit cube, not on a subset of the sphere. A Lipschitz constant of $1(\text{m}^{2}/\text{s}^{2})/\text{km}$ is used for selecting the number of neighbors.
E.7 Model Selection on Synthetic Data
Figure 13: Training (blue), validation (orange), and test (green) data for a single seed of the model selection experiment. The dashed red line depicts predictive method $h_{0}$, and dashed black shows $h_{1}$.
We next see that SNN and 1NN are able to select the model with lower test risk in a model selection task, but the holdout systematically picks the wrong model. We repeat the model selection problem 100 times. In each repetition, we have $N^{\text{train}}=100$ and a max $N^{\textup{{val}}}=75$. In our analysis, we will consider validation subsets of size $N^{\textup{{val}}}\in\{5\ell\}_{\ell=1}^{15}$. We generate independent test, validation, and training data as follows; see Figure 13.
@@ E.7 Model Selection on Synthetic Data
We compare two predictive methods: $h_{0}(S)=0.25$ and $h_{1}(S)=\beta_{1}^{\top}S+\beta_{0}$, with $(\beta_{1},\beta_{0})$ fit by minimizing the mean absolute residual on the training data. Figure 13 shows the data and predictions of both models (as functions of space). We use the loss function $\ell(a,b)=|a-b|$, which is bounded for this problem because both the hypotheses and the response variable are bounded on $[0,1]$.
Across all seeds, $h_{0}$ has the lower empirical test risk; $h_{1}$ makes large errors on the test points near 00 because most of the training data is in $[0,0.5]$. Since most of the validation data also clusters near 1, we expect the holdout to struggle due to bias. Our arguments in Appendix C lead us to expect both SNN and 1NN should perform well on this task when given sufficient validation data.
We say an estimator of the risk, $\hat{R}$, selects $h_{0}$ if $\hat{R}(h_{0})<\hat{R}(h_{1})$. We plot the percentage of times each method (correctly) selects $h_{0}$ as a function of the number of validation points in Figure 14. When the validation set is small, all estimators select the model with lowest test risk ($h_{0}$) less than half the time. For the nearest neighbor methods, we expect that when there are few or no spatial locations less than $1$, weighting cannot fix the estimate. However, when the number of validation points is large, the nearest neighbor methods consistently (correctly) select $h_{0}$. By contrast, the holdout consistently (incorrectly) selects $h_{1}$, even though $h_{1}$ has higher test risk. See Appendix C for full experiment details.
@@
The data generation is fully described by Equation 137.
@@
We consider two models. The first is a constant predictor that predicts $0.25$. The second is an affine model (a linear model with an intercept) fit by minimizing the mean absolute error from the line to the training points. This is fit using the Scikit-learn quantile regression with the (default) “HiGHS” solver [Huangfu and Hall, 2015].
@@
The validation estimates used are calculated in the same as the synthetic experiments outlined previously in Section E.3.4. We use $\Delta=1$ in the bound when selecting $k^{\star}_{T_{2}}$, even though the absolute value loss used can be larger than $1$. We don’t expect this to have a significant impact on the results, as the upper bound we minimize is already misspecified in a similar way by not using the actual Lipschitz constant of the function. We again use $\delta=0.1$ when selecting $k^{\star}_{T_{2}}$.

### v3

@@ D.6.2 Loss Function
D.6.3 Estimation of Ground Truth
Following the argument in section D.1, we use the empirical test risk eq. 114 in place of the test risk as ground truth in synthetic experiments. The assumption that $(Z_{m})_{m=1}^{M^{\textup{{test}}}}$ are independent. It is sufficient for $Y^{\textup{{test}}}_{m}$ (the wind daily wind speeds) to be independent, conditioned on the spatial location at which it is observed. This is likely not the case, as we would expect average wind speed in consecutive days exhibit at least some dependence. However, if the wind speed decorrelates reasonably rapidly over time, we would expect similar arguments to hold, possibly with fewer effective samples.
Because the loss is truncated mean squared error, the $Z_{m}$ are surely bounded by $25$m2/s2. We use $775$ samples in estimating the risk. Using these numbers, and under the assumption that wind speed at a location is independent of the wind speed on previous days, in eq. 117, we arrive at
@@ D.6.3 Estimation of Ground Truth
Comparing to fig. 15, we see that this application of Hoeffding’s inequality is not sufficient to justify that the estimate of ground truth is accurate enough to attributed the observed better performance of SNN to (actually) better estimation of the ground truth as opposed to inaccuracies of our Monte Carlo estimate of the test risk. However, we expect this is largely due to looseness is Hoeffding’s inequality and, given that a large difference is observed in most seeds, it would be very surprising if this was only due to error in estimation of the ground truth which is independent across seeds.
D.6.4 Model Fitting
A gradient boosted machine is fit using LightGBM (Ke et al., 2017) with default parameters expect for the number of leaves (set to 127) and the number of estimators (set to 100).
D.6.5 Risk Estimation procedures
Holdout, 1NN and SNN are run as in the air temperature experiments (section D.4.5). In particular, we use a failure probability of $\delta=0.1$ for SNN and nearest neighbor calculations are done with respect to Haversine distance to account for Earth’s curvature. We use a fixed $\delta$ since we only discuss a method applicable to estimating fill distance on the unit cube, not on a subset of the sphere. A Lipschitz constant of $1(\text{m}^{2}/\text{s}^{2})/\text{km}$ is used for selecting the number of neighbors.
D.7 Model Selection on Synthetic Data
Figure 16: Training (pink circles), validation (orange diamonds), and test (blue triangles) data for a single seed of the model selection experiment. The dashed red line depicts predictive method $h_{0}$, and dashed black shows $h_{1}$.
We next see that SNN and 1NN are able to select the model with lower test risk in a model selection task, but the holdout systematically picks the wrong model. We repeat the model selection problem 100 times. In each repetition, we have $N^{\text{train}}=100$ and a max $N^{\textup{{val}}}=75$. In our analysis, we will consider validation subsets of size $N^{\textup{{val}}}\in\{5\ell\}_{\ell=1}^{15}$. We generate independent test, validation, and training data as follows; see fig. 16.
@@ D.7 Model Selection on Synthetic Data
We compare two predictive methods: $h_{0}(S)=0.25$ and $h_{1}(S)=\beta_{1}^{\top}S+\beta_{0}$, with $(\beta_{1},\beta_{0})$ fit by minimizing the mean absolute residual on the training data. fig. 16 shows the data and predictions of both models (as functions of space). We use the loss function $\ell(a,b)=|a-b|$, which is bounded for this problem because both the hypotheses and the response variable are bounded on $[0,1]$.
Across all seeds, $h_{0}$ has the lower empirical test risk; $h_{1}$ makes large errors on the test points near 00 because most of the training data is in $[0,0.5]$. Since most of the validation data also clusters near 1, we expect the holdout to struggle due to bias. Our arguments in appendix B lead us to expect both SNN and 1NN should perform well on this task when given sufficient validation data.
We say an estimator of the risk, $\hat{R}$, selects $h_{0}$ if $\hat{R}(h_{0})<\hat{R}(h_{1})$. We plot the percentage of times each method (correctly) selects $h_{0}$ as a function of the number of validation points in fig. 17. When the validation set is small, all estimators select the model with lowest test risk ($h_{0}$) less than half the time. For the nearest neighbor methods, we expect that when there are few or no spatial locations less than $1$, weighting cannot fix the estimate. However, when the number of validation points is large, the nearest neighbor methods consistently (correctly) select $h_{0}$. By contrast, the holdout consistently (incorrectly) selects $h_{1}$, even though $h_{1}$ has higher test risk. See appendix B for full experiment details.
@@
The data generation is fully described by eq. 139.
@@
We consider two models. The first is a constant predictor that predicts $0.25$. The second is an affine model (a linear model with an intercept) fit by minimizing the mean absolute error from the line to the training points. This is fit using the Scikit-learn quantile regression with the (default) “HiGHS” solver (Huangfu and Hall, 2015).
@@
The validation estimates used are calculated in the same as the synthetic experiments outlined previously in section D.3.4. We use $\Delta=1$ in the bound when selecting $k^{\star}_{T_{2}}$, even though the absolute value loss used can be larger than $1$. We don’t expect this to have a significant impact on the results, as the upper bound we minimize is already misspecified in a similar way by not using the actual Lipschitz constant of the function. We again use $\delta=0.1$ when selecting $k^{\star}_{T_{2}}$.